# CaPo: Cooperative Plan Optimization for Efficient Embodied Multi-Agent Cooperation

### v1

@@ Abstract
In this work, we address the cooperation problem among large language model (LLM) based embodied agents, where agents must cooperate to achieve a common goal. Previous methods often execute actions extemporaneously and incoherently, without long-term strategic and cooperative planning, leading to redundant steps, failures, and even serious repercussions in complex tasks like search-and-rescue missions where discussion and cooperative plan are crucial. To solve this issue, we propose Cooperative Plan Optimization (CaPo) to enhance the cooperation efficiency of LLM-based embodied agents. Inspired by human cooperation schemes, CaPo improves cooperation efficiency with two phases: 1) meta-plan generation, and 2) progress-adaptive meta-plan and execution. In the first phase, all agents analyze the task, discuss, and cooperatively create a meta-plan that decomposes the task into subtasks with detailed steps, ensuring a long-term strategic and coherent plan for efficient coordination. In the second phase, agents execute tasks according to the meta-plan and dynamically adjust it based on their latest progress (e.g., discovering a target object) through multi-turn discussions. This progress-based adaptation eliminates redundant actions, improving the overall cooperation efficiency of agents. Experimental results on the ThreeDworld Multi-Agent Transport and Communicative Watch-And-Help tasks demonstrate CaPo’s much higher task completion rate and efficiency compared with state-of-the-arts.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and generating human language, complex reasoning, and planning, achieving impressive performance (OpenAI, 2024; Touvron et al., 2023). These advancements empower LLM-based embodied agents to autonomously make plans (Li et al., 2023a; Padmakumar et al., 2022; Zhu et al., 2023; Wang et al., 2023; Wu et al., 2023b; Huang et al., 2022b) and perform reasoning (Du et al., 2023; Hao et al., 2023; Zhou et al., 2024; Huang et al., 2022a) by using human language to assist people in daily activities, such as housework and daily chores. The next milestone for agents is to cooperate with others to achieve joint tasks. This is crucial not only for efficiently performing simple tasks but also for tackling complex ones that cannot be completed in isolation due to their inherent complexity or the dynamic nature of the environment (Zhang et al., 2023b; Guo et al., 2024; Mandi et al., 2023; Zhang et al., 2023a).
@@
Figure 1: Procedure example of task accomplishment of CoELA (Zhang et al., 2023b) and our CaPo. In CoELA, after each action execution, Alice and Bob communicate to decide next action which is a greedy single-step plan and suboptimal. For example, they do not use wood basket which can contain sever objects, and both extemporaneously move a single item to the target bed without a long-term strategic and collaborative plan. Differently, in CaPo, Alice and Bob first discuss to make a long-term meta-plan for strategical cooperation in which Alice is arranged to move several target items into a wood basket, and Bob moves the remaining target items and also searches the unknown objects. Then during execution phase, both follow the meta-plan to accomplish task, and dynamically adapt the meta-plan the latest task progress, ensuring its effectiveness and efficiency in coordinating agents.
@@
Notably, the cooperation among LLM-based embodied agents is rarely investigated despite being highly desired. Conventional works often focus on adopting reinforcement learning (RL) (Jiang & Lu, 2018; Liu et al., 2021; Wang et al., 2021) to explore the dynamics of cooperative behavior among non-LLM-based agents. In spite of their promising performance in certain scenarios, RL-based cooperation methods exhibit limited adaptability across different tasks (Dittadi et al., 2021; Cobbe et al., 2019), since they are often not trained on large-scale data and lack sufficient generalization ability. To solve this issue, in this work, we are particularly interested in the problem of “how to develop an effective collaboration framework for LLM-based agents”, since LLMs have revealed strong reasoning, planning, and communication ability across different tasks and thus are regarded as good agents’ brains.
@@
Among the limited related works, CoELA (Zhang et al., 2023b) proposes an LLM-based multi-agent cooperation framework in which after each action execution, agents communicate to devise a single-step plan for the next action. Despite its significant advancements, CoELA’s short-term, single-step planning, which lacks consideration for long-term strategic collaboration, often results in extemporaneous and incoherent actions among agents, leading to several potential issues. Firstly, without a long-term coherent collaboration plan, it leads to numerous redundant action steps and increased costs, since agents’ movement is not easy and is indeed expensive in the physical world. For instance, as shown in Fig. 1, for the object transport task, agent Alice and Bob do not use the wood basket which can contain several objects, and extemporaneously move their nearest target objects one by one, leading to inferior efficiency. Moreover, complex tasks are difficult to accomplish without thorough discussion and long-term collaboration, especially in (embodied) environments where each agent has only partial observations. Finally, without a long-term cooperative plan, agents’ extemporaneous actions can result in mistakes with severe consequences. For instance, in search-and-rescue missions, poor coordination can have dire outcomes, such as endangering human lives due to the complex nature of these operations.
@@
Contributions. To address the above issues, we propose a novel and effective Cooperative Plan Optimization (CaPo) framework that uses LLMs’ strong reasoning and planning ability to enhance the cooperation efficiency of LLM-based embodied agents. Inspired by human cooperation schemes (Tuomela, 1998; Thürmer et al., 2017), CaPo engages agents in multi-turn discussions to create and update a long-term strategic and coherent meta-plan, providing step-by-step guidance to coordinate agents and efficiently complete tasks.
Specifically, to accomplish a task, CaPo consists of two phases: 1) meta-plan generation, providing long-term strategical and coherent guidance for coordinating agents, and 2) progress-adaptive meta-plan and execution, dynamically adapting the meta-plan to agents’ latest progress. In the first phase, agents analyze the task and discuss with other agents for collecting relevant information. Next, one agent is responsible for making a meta-plan which decomposes the task into subtasks with detailed accomplishment steps like agent allocations, and then collects the feedback from other agents for further meta-plan refinement. The steps of meta-plan generation and refinement will continue until all agents reach a consensus or the communication cost is exhausted. This approach ensures the thorough discussion and analysis of all agents, helping to make a long-term strategical and coherent meta-plan for efficiently coordinating all agents. For example, as illustrated in Fig. 1, in the object transport task, agents Alice and Bob are strategically assigned to different subtasks.
In the second phase, as shown in Fig. 1, agents follow the meta-plan from the first phase, and focus on their assigned subtasks. As progress is made, agents may complete subtasks or make important observations, such as Alice in Fig. 1 discovering the object “iPhone” which is Bob’s target. Accordingly, agents dynamically adapt meta-plan to the latest task progress through multi-turn discussions, allowing Alice to handle the object “iPhone” and complete the task efficiently. This progress-adaptive approach ensures that the meta-plan remains effective in coordinating all agents, thereby enhancing cooperation efficiency.
@@
Finally, experimental results demonstrate that CaPo significantly improves task completion rates and efficiency compared to state-of-the-art (SoTA) methods on the widely used ThreeDworld Multi-Agent Transport task (Zhang et al., 2023b) (object transport task) and the Communicative Watch-And-Help task (Zhang et al., 2023b) (household chore task). For instance, on the ThreeDworld Multi-Agent Transport task, CaPo surpasses the SoTA CoELA by 16.7% and 4.7% in completion rate with GPT-3.5 and GPT-4 based agents, respectively
@@ 2 Related Work
LLM-based Agents. LLM-based agents (Hong et al., 2023; Wang et al., 2024; Shen et al., 2024; Liu et al., 2023a) are designed to autonomously perceive environments, execute actions, accumulate knowledge, and evolve themselves, with rich real world knowledge and complex reasoning capability inherited from LLMs. Notable agents like AutoGPT (Richards & et al, 2021), BabyAGI (Nakajima, 2023), and AgentGPT (Reworkd, 2023) showcase remarkable proficiency in decision-making and complex reasoning. In the embodied environment, LLM-based agents have shown superior capacity in strategic planning (Li et al., 2023a; Padmakumar et al., 2022; Wu et al., 2023b; Huang et al., 2022b). Specifically, LLM-planner (Song et al., 2023) harness LLMs to do few-shot planning for embodied agents. PET (Wu et al., 2023a) translates a task description with LLMs into a list of high-level sub-tasks. TaPA wu2023embodied enables the agent to generate executable plans by aligning LLMs with visual perception models. Another line of research focuses on harnessing LLMs’s reasoning capabilities in embodied tasks (Zhou et al., 2024; Huang et al., 2022a). ELLM (Du et al., 2023) utilizes LLMs to set pretraining goals in RL, guiding agents towards the goal without human involvement.
@@ 3 Preliminaries
We follow previous work (Zhang et al., 2023b; Gong et al., 2023) and formulate the embodied multi-agent cooperation task as an decentralized partially observable Markov decision process (DEC-POMDP) (Bernstein et al., 2002; Spaan et al., 2006), which is defined as $<n,\mathcal{S},\mathcal{O},\mathcal{A},P,r,\gamma>$. Here, $n$ represents the number of agents; $\mathcal{S}$ is the finite state space; $\mathcal{O}$ denotes the observation space; $\mathcal{A}$ is a finite joint action space of all agents; $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability function; $r=\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function; $\gamma\in[0,1]$ denotes the discount factor. In this framework, at time step $t\in\mathbb{N}$, each agent $i$ observes the environment’s state $s_{t}\in\mathcal{S}$, and receives an observation set $\mathcal{O}_{i}$. $\mathcal{O}_{i}$ consists of a world observation $\mathcal{O}_{i}^{w}$, which the agent gathers through its sensors, or a communication message observation $\mathcal{O}_{i}^{c}$ from other teammate agents. Agent $i$ takes actions from its action space $\mathcal{A}_{i}$, which includes a finite set of world action $\mathcal{A}_{i}^{w}$, e.g., grasping a target object, or a finite set of messaging action $\mathcal{A}_{i}^{c}$. Then agents receive a shared reward $r_{t}=r(s_{t},a_{t})$, where $a_{t}\in\mathcal{A}$ denotes the joint actions of agents, and observe a new state $s_{t+1}$ with probability $P(s_{t+1}|s_{t},a_{t})$. We formulate the problem with two decentralized intelligent embodied agents working together to complete a long-horizon rearrangement task (Zhang et al., 2023b; Batra et al., 2020) in a multi-room indoor environment. During the task, agents can execute multiple kinds of actions, such as navigation, interaction, and communication by sending messages.
@@ 4.1 Overall Framework of CaPo
CaPo aims to enhance cooperation efficiency of LLM-based embodied agents. Its key idea is to create a long-term meta-plan for strategically and coherently coordinating agents to complete a rearragement task. Accordingly, agents follow the meta-plan to complete task step by step, and dynamically adapt the meta-plan to their latest progress, thereby avoiding redundant work allocation and improving overall cooperation efficiency.
The overall pipeline of CaPo is in Fig. 2. Each agent contains several modules, including 1) a perception module, 2) a memory module, 3) a communication module, 4) a cooperative planning module, 5) a progress-adaptive planning module, 6) a plan parsing module, and 7) an execution module. For each agent, the perception module gathers observations from environment, including messages from other agents and relevant scene information from the RGB-D image. The memory module dynamically stores the shared task, dialogue history between agents, agent progress, teammate progress, and action history, all formatted as text descriptions. Additionally, the semantic map from the perception module is also stored. The communication module retrieves relevant information from the memory module and uses an LLM to generate messages that are sent to other agents. The cooperative planning module either generates the meta-plan or provides feedback on the meta-plan, as seen with Alice and Bob in Fig. 2, respectively. The progress-adaptive planning module adapts the meta-plan to agents’ latest task progress. The plan parsing module, powered by the LLM, determines which sub-plan to execute based on relevant information from the memory module and available actions for the current state. Finally, the execution module converts high-level sub-plans from the plan parsing module into primitive actions for execution.
@@ 5 Experiments
Metrics. On TDW-MAT, we adopt Transport Rate, i.e., the fraction of subtasks completed within 3000 time steps (a.k.a. frames), as performance metric. Note, one action step may last multiple time steps, e.g., resetting arms. On C-WAH, Average Steps to complete all tasks is used as the metric to evaluate cooperation efficiency.
Implementation. Following CoELA, we test two settings on TDW-MAT task: 1) a real-world setting where the perception module is instantiated as Mask-RCNN (He et al., 2017) that is trained using collected scene images (Zhang et al., 2023b), and 2) an oracle setting with segmentation ground-truth. We use GPT-3.5-turbo and GPT-4 from the OpenAI API (OpenAI, 2024), and LLAMA-2-13B-CHAT (Touvron et al., 2023), as LLMs in embodied agents. We set default parameters for LLMs: temperature of 0.7, a maximum of 256 output tokens, and top-p sampling with $p=1$. Our code will be made publicly available.
Classic Agents GPT-3.5 Agents LLAMA-2 Agents
RHP RHP CoELA ProAgent RoCo CaPo${}_{\text{(ours)}}$ CoELA ProAgent RoCo CaPo${}_{\text{(ours)}}$
@@
[…]
Stuff ($\uparrow$) 36 54 39 44+12.8% 47+20.1% 45+15.4% 48 55${}_{\ +14.6\%}$ 51${}_{\ +6.3\%}$ 56${}_{\ +16.7\%}$
Avg. ($\uparrow$) 43 61 52 56${}_{\ +7.7\%}$ 55${}_{\ +5.6\%}$ 57${}_{\ +9.6\%}$ 53 ${}_{\ +7.5\%}$ 55${}_{\ +3.8\%}$ 61${}_{\ +15.1\%}$
@@
Food ($\uparrow$) 52 76 72 80${}_{\ +11.1\%}$ 74${}_{\ +2.8\%}$ 85${}_{\ +18.1\%}$ 60 64${}_{\ +6.7\%}$ 62${}_{\ +3.3\%}$ 66${}_{\ +10.0\%}$
[…]
Avg. ($\uparrow$) 50 75 72 78${}_{\ +8.3\%}$ 77${}_{\ +6.9\%}$ 84${}_{\ +16.7\%}$ 62 63${}_{\ +1.6\%}$ 65${}_{\ +4.8\%}$ 71${}_{\ +14.5\%}$
Table 1: Comparison of average Transport Rate (TR, %) of all baselines on the TDW-MAT w/o and w/ Oracle Perception task. Each task requires agents to move two kinds of items, including Food and Stuff. RHP uses a single agent while all others adopt two agents. The subscript value like ${+9.6\%}$ in 57${}_{\ +9.6\%}$ denotes the relative improvement when comparing with baseline CoELA.
Classic Agents GPT-4 Agents
MHP MHP CoELA ProAgent RoCo CaPo${}_{\text{(ours)}}$
[…]
Visual Obs ($\downarrow$) 141 103 92 90${}_{\ +2.2\%}$ 89${}_{\ +3.2\%}$ 83${}_{\ +9.8\%}$
Table 2: Comparison of Average Step (AS) of all compared methods on the C-WAH task. Symbolic obs and visual obs denote symbolic observation and visual observation settings, respectively. MHP uses a single agent while all others adopt two agents.
Baselines. We adopt two types of methods as our baseline: 1) classical agents, including MCTS-based Hierarchical Planner (MHP(Puig et al., 2020) and Rule-based Hierarchical Planner (RHP(Gan et al., 2022). 2) LLM-driven agents, including CoELA Zhang et al. (2023b), ProAgent (Zhang et al., 2023a), and RoCo (Mandi et al., 2023). CoELA Zhang et al. (2023b) features a modular framework for multi-agent planning, communication, and complete long-horizon tasks, but generate independent short-term plan for each agent. ProAgent (Zhang et al., 2023a), and RoCo (Mandi et al., 2023) generate joint plans for cooperative agents, and introduce a reflection loop or environment feedback for plan validation. See more details in Appendix A.1.

### v2

@@ Abstract
In this work, we address the cooperation problem among large language model (LLM) based embodied agents, where agents must cooperate to achieve a common goal. Previous methods often execute actions extemporaneously and incoherently, without long-term strategic and cooperative planning, leading to redundant steps, failures, and even serious repercussions in complex tasks like search-and-rescue missions where discussion and cooperative plan are crucial. To solve this issue, we propose Cooperative Plan Optimization (CaPo) to enhance the cooperation efficiency of LLM-based embodied agents. Inspired by human cooperation schemes, CaPo improves cooperation efficiency with two phases: 1) meta-plan generation, and 2) progress-adaptive meta-plan and execution. In the first phase, all agents analyze the task, discuss, and cooperatively create a meta-plan that decomposes the task into subtasks with detailed steps, ensuring a long-term strategic and coherent plan for efficient coordination. In the second phase, agents execute tasks according to the meta-plan and dynamically adjust it based on their latest progress (e.g., discovering a target object) through multi-turn discussions. This progress-based adaptation eliminates redundant actions, improving the overall cooperation efficiency of agents. Experimental results on the ThreeDworld Multi-Agent Transport and Communicative Watch-And-Help tasks demonstrate CaPo’s much higher task completion rate and efficiency compared with state-of-the-arts. The code is released at https://github.com/jliu4ai/CaPo.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and generating human language, complex reasoning, and planning, achieving impressive advancement (OpenAI, 2024; Touvron et al., 2023). These advancements empower LLM-based embodied agents to autonomously make plans (Li et al., 2023a; Padmakumar et al., 2022; Zhu et al., 2023; Wang et al., 2023a; Wu et al., 2023b; Huang et al., 2022b) and perform reasoning (Du et al., 2023; Hao et al., 2023; Zhou et al., 2024; Huang et al., 2022a) by using human language to assist people in daily activities, such as housework and daily chores. The next milestone for agents is to cooperate with others to achieve joint tasks. This is crucial not only for efficiently performing simple tasks but also for tackling complex ones that cannot be completed in isolation due to their inherent complexity or the dynamic nature of the environment (Zhang et al., 2023b; Guo et al., 2024; Mandi et al., 2023).
@@
Figure 1: Examples of task accomplishment of CoELA (Zhang et al., 2023b) and our CaPo. In CoELA, after each action execution, Alice and Bob communicate to decide next action which is a greedy single-step plan and suboptimal. For example, they do not use wood basket which can contain multiple objects, and both extemporaneously move a single item to the target bed without a long-term and collaborative plan. In contrast, in CaPo, Alice and Bob first discuss to make a long-term meta-plan for strategical cooperation in which Alice is arranged to move several target items into a wood basket, and Bob moves the remaining target items and also searches the unknown objects. Then during execution phase, both follow the meta-plan to accomplish task, and dynamically adapt the meta-plan to latest task progress, ensuring its effectiveness and efficiency in coordinating agents.
@@
Notably, the cooperation among LLM-based embodied agents is rarely investigated despite being highly desired. Conventional works often focus on adopting reinforcement learning (RL) (Jiang & Lu, 2018; Liu et al., 2021; Wang et al., 2021) to explore the dynamics of cooperative behavior among non-LLM-based agents. In spite of their promising performance in certain scenarios, RL-based cooperation methods exhibit limited adaptability across different tasks (Dittadi et al., 2021; Cobbe et al., 2019), since they are often not trained on large-scale data and lack sufficient generalization ability. To solve this issue, in this work, we are particularly interested in the problem of “how to develop an effective collaboration framework for LLM-based agents”, since LLMs have revealed strong reasoning and planning capability as good agents’ brain across different tasks.
@@
Among the limited related works, CoELA (Zhang et al., 2023b) proposes an LLM-based multi-agent cooperation framework in which after each action execution, agents communicate to devise a single-step plan for the next action. Despite its significant advancements, CoELA’s short-term, single-step planning, which lacks consideration for long-term strategic collaboration, often results in extemporaneous and incoherent actions among agents, leading to several potential issues. ① Without a long-term coherent collaboration plan, agents are prone to executing numerous redundant actions, resulting in increased costs. This issue is particularly critical in the physical world, where agents’ movements are inherently challenging and resource-intensive. For instance, as shown in Fig. 1, for the object transport task, agent Alice and Bob do not use the wood basket which can contain several objects, and extemporaneously move their nearest target objects one by one, leading to inferior efficiency. ② Complex tasks are challenging to accomplish without thorough discussion and long-term collaboration, particularly in embodied environments where agents operate with only partial observations. ③ Without a long-term cooperative plan, agents’ uncoordinated actions can lead to severe consequences. For example, in search-and-rescue missions, poor coordination may endanger lives due to the complexity of such operations.
@@
Contributions. This work introduces Cooperative Plan Optimization (CaPo), a novel framework designed to enhance the cooperation efficiency of LLM-based embodied agents by leveraging their advanced reasoning and planning capabilities. Inspired by human cooperation schemes (Tuomela, 1998; Thürmer et al., 2017), CaPo enables agents to collaboratively create and refine long-term, strategic, and coherent meta-plans through multi-turn discussions. These meta-plans serve as actionable guides that facilitate efficient task coordination and completion.
Specifically, CaPo addresses key challenges in multi-agent cooperation through two main phases. The first phase is Meta-Plan Generation, where agents collaboratively analyze the task and exchange information to generate a meta-plan that decomposes the overall task into detailed subtasks, including agent assignments and execution steps. A designated agent initiates the meta-plan, while others evaluate and refine it through iterative feedback, ensuring alignment and optimal planning. The discussion process continues until consensus is reached or the communication budget is exhausted. This phase ensures thorough deliberation and creates a coherent roadmap for task execution. For example, in the object transport task, agents Alice and Bob are strategically assigned complementary subtasks to optimize their coordination (see Fig. 1).
The second phase is Progress-Adaptive Meta-Plan and Execution, where agents dynamically adapt the meta-plan based on real-time progress and discoveries during task execution. For instance, if Alice identifies the location of an object that was initially unknown (e.g., an “iPhone”), the agents collaboratively update the meta-plan to incorporate this new information, ensuring timely and efficient task completion. This adaptive mechanism maintains the meta-plan’s effectiveness, even in dynamic environments, and prevents inefficiencies or redundant actions.
@@
The proposed approach has been rigorously evaluated on benchmark tasks, including the ThreeDworld Multi-Agent Transport (TDW-MAT) and Communicative Watch-And-Help (C-WAH) tasks (Zhang et al., 2023b). Experimental results demonstrate that CaPo significantly improves task completion rates and cooperation efficiency compared to state-of-the-art methods. Notably, on the TDW-MAT task, CaPo achieves a completion rate improvement of 16.7% and 4.7% over the SoTA CoELA method with GPT-3.5 and GPT-4 agents, respectively.
@@ 2 Related Work
LLM-based Agents. LLM-based agents (Hong et al., 2023; Wang et al., 2024; Shen et al., 2024; Liu et al., 2023a) are designed to autonomously perceive environments, execute actions, and accumulate knowledge, with rich real world knowledge and complex reasoning capability inherited from LLMs. Notable agents like AutoGPT (Richards & et al, 2021), BabyAGI (Nakajima, 2023), and AgentGPT (Reworkd, 2023) showcase remarkable proficiency in decision-making and complex reasoning. In the embodied environment, LLM-based agents have shown superior capacity in strategic planning (Li et al., 2023a; Padmakumar et al., 2022; Wu et al., 2023b; Huang et al., 2022b). Specifically, LLM-planner (Song et al., 2023) harness LLMs to do few-shot planning for embodied agents. PET (Wu et al., 2023a) translates a task description with LLMs into a list of high-level sub-tasks. TaPA wu2023embodied enables the agent to generate executable plans by aligning LLMs with visual perception models. Another research line focuses on harnessing LLMs’s reasoning capabilities in embodied tasks (Zhou et al., 2024; Huang et al., 2022a). ELLM (Du et al., 2023) utilizes LLMs to set pretraining goals in RL, guiding agents towards the goal without human involvement.
@@ 3 Preliminaries
Task Formulation. We formulate the embodied multi-agent cooperation task as an decentralized partially observable Markov decision process (DEC-POMDP) (Bernstein et al., 2002; Spaan et al., 2006), which is defined as $<n,\mathcal{S},\mathcal{O},\mathcal{A},P,r,\gamma>$. Here, $n$ represents the number of agents; $\mathcal{S}$ is the finite state space; $\mathcal{O}$ denotes the observation space; $\mathcal{A}$ is a finite joint action space of all agents; $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability function; $r=\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function; $\gamma\in[0,1]$ denotes the discount factor. In this framework, at time step $t\in\mathbb{N}$, each agent $i$ observes the environment’s state $s_{t}\in\mathcal{S}$, and receives an observation set $\mathcal{O}_{i}$. $\mathcal{O}_{i}$ consists of a world observation $\mathcal{O}_{i}^{w}$, which the agent gathers through its sensors, or a communication message observation $\mathcal{O}_{i}^{c}$ from other teammate agents. Agent $i$ takes actions from its action space $\mathcal{A}_{i}$, which includes a finite set of world action $\mathcal{A}_{i}^{w}$, e.g., grasping a target object, or a finite set of messaging action $\mathcal{A}_{i}^{c}$. Then agents receive a shared reward $r_{t}=r(s_{t},a_{t})$, where $a_{t}\in\mathcal{A}$ denotes the joint actions of agents, and observe a new state $s_{t+1}$ with probability $P(s_{t+1}|s_{t},a_{t})$. We formulate the problem with two decentralized intelligent embodied agents working together to complete a long-horizon rearrangement task (Zhang et al., 2023b; Batra et al., 2020) in a multi-room indoor environment. During the task, agents can execute multiple kinds of actions, such as navigation, interaction, and communication by sending messages.
CoELA Framework. CoELA Zhang et al. (2023b) is a pioneering modular framework for embodied multi-agent cooperation, which consists of five key modules: (a) Perception, (b) Memory, (c) Communication, (d) Planning, and (e) Execution. For each agent, the (a) Perception Module gathers observations from environment, including messages from other agents and relevant scene information from the RGB-D image. The (b) Memory Module dynamically stores the shared task, dialogue history between agents, agent progress, team- mate progress, and action history, all formatted as text descriptions. (c) The Communication Module retrieves relevant information from the memory module and uses an LLM to generate messages that are sent to other agents. (d) The Planning Module driven by LLM decides which plan to take given related information retrieved from the memory module and available actions proposed regarding the current state. Finally, the (e) Execution Module convert high-level plan into primitive actions executable in the environment. CoELA exhibits great performances in the embodied multi-agent cooperation tasks, making it a strong baseline for validating the effectiveness of our model in improving multi-agent cooperation efficiency.
@@ 4.1 Overall Framework of CaPo
CaPo addresses the challenge of enabling two centralized embodied agents to effectively cooperate in a shared environment. In this setup, each agent has partial observations and must rely on communication and coordination to accomplish complex tasks. The key objective is to achieve strategic and step-by-step collaboration, where both agents contribute to the task’s completion through efficient planning and adaptive decision-making. As shown in Fig. 2, each agent is equipped with five foundational modules from CoELA: perception, memory, communication, plan-phrasing, and execution. These modules enable the agents to perceive their environment, store and retrieve relevant information, exchange messages for coordination, generate plans, and execute actions accordingly.
@@ 5 Experiments
Metrics. On TDW-MAT, we adopt Transport Rate, i.e., the fraction of subtasks completed within 3000 time steps (a.k.a. frames), as performance metric. Note, one action step may last multiple time steps, e.g., resetting arms. On C-WAH, Average Steps to complete all tasks is used as the metric to evaluate cooperation efficiency. Following CoELA (Zhang et al., 2023b), we also report Efficiency Improvement (EI) of cooperating with other agents as $\Delta M/M_{0}$, where $\Delta M$ is the main efficiency metric difference, and $M_{0}$ denotes the larger on of the main efficiency metric for numerical stability.
Implementation. We test two settings on TDW-MAT task: 1) a real-world setting where the perception module is instantiated as Mask-RCNN (He et al., 2017) that is trained using collected scene images (Zhang et al., 2023b), and 2) an oracle setting with segmentation ground-truth. We use GPT-3.5-turbo and GPT-4 from the OpenAI API (OpenAI, 2024), and LLAMA-2-13B-CHAT (Touvron et al., 2023), as LLMs in embodied agents. We set default parameters for LLMs: temperature of 0.7, a maximum of 256 output tokens, and top-1 sampling. Our code will be made publicly available.
Classic Agents GPT-3.5 Agents LLAMA-2 Agents GPT-4 Agents
RHP∗† RHP CoELA CaPo${}_{\text{(ours)}}$ CoELA CaPo${}_{\text{(ours)}}$ CoELA ProAgent RoCo CaPo${}_{\text{(ours)}}$
@@
Food ($\uparrow$) 49 67${}_{\ +25\%}$ 67${}_{\ +23\%}$ 70${}_{\ +31\%}$ 57${}_{\ +9\%}$ 66${}_{\ +17\%}$ 82${}_{\ +38\%}$ 82${}_{\ +27\%}$ 83 34% 85${}_{\ +43\%}$
Stuff ($\uparrow$) 36 54${}_{\ +34\%}$ 39${}_{\ +18\%}$ 45${}_{\ +27\%}$ 48${}_{\ +11\%}$ 56${}_{\ +22\%}$ 61${}_{\ +41\%}$ 57${}_{\ +35\%}$ 60${}_{\ +39\%}$ 64${}_{\ +40\%}$
Avg. ($\uparrow$) 43 61${}_{\ +29\%}$ 52${}_{\ +20\%}$ 57${}_{\ +29\%}$ 53${}_{\ +10\%}$ 61${}_{\ +19\%}$ 71${}_{\ +39\%}$ 69${}_{\ +31\%}$ 71${}_{\ +36\%}$ 74${}_{\ +41\%}$
@@
Food ($\uparrow$) 52 76${}_{\ +33\%}$ 72${}_{\ +29\%}$ 85${}_{\ +38\%}$ 60${}_{\ +3\%}$ 66${}_{\ +14\%}$ 87${}_{\ +41\%}$ 84${}_{\ +37\%}$ 88${}_{\ +42\%}$ 90${}_{\ +40\%}$
Stuff ($\uparrow$) 49 74${}_{\ +34\%}$ 73${}_{\ +32\%}$ 84${}_{\ +39\%}$ 63${}_{\ +19\%}$ 76${}_{\ +23\%}$ 83${}_{\ +41\%}$ 85${}_{\ +34\%}$ 82${}_{\ +35\%}$ 87${}_{\ +38\%}$
Avg. ($\uparrow$) 50 75${}_{\ +34\%}$ 72${}_{\ +30\%}$ 84${}_{\ +38\%}$ 62${}_{\ +8\%}$ 71${}_{\ +18\%}$ 85${}_{\ +41\%}$ 84${}_{\ +35\%}$ 85${}_{\ +38\%}$ 89${}_{\ +39\%}$
Table 1: Comparison of average Transport Rate(%) of all baselines on the TDW-MAT w/o and w/ Oracle Perception task. Each task requires agents to move two kinds of items, including Food and Stuff. RHP uses a single agent while all others adopt two agents. denotes results quoted from CoELA. The subscript value like ${+25\%}$ in 67${}_{\ +25\%}$ denotes the Efficiency Improvement.
Classic Agents Heterogeneous Agents GPT-4 Agents
MHP∗† MHP MHP+CoELA MHP+CaPo CoELA ProAgent RoCo CaPo${}_{\text{(ours)}}$
Symbolic Obs ($\downarrow$) 111 75${}_{\ +33\%}$ 59${}_{\ +45\%}$ 57${}_{\ +47\%}$ 57${}_{\ +49\%}$ 62${}_{\ +37\%}$ 57${}_{\ +43\%}$ 51${}_{\ +46\%}$
Visual Obs ($\downarrow$) 141 103${}_{\ +26\%}$ 94${}_{\ +34\%}$ 90${}_{\ +38\%}$ 92${}_{\ +34\%}$ 90${}_{\ +37\%}$ 89${}_{\ +32\%}$ 83${}_{\ +37\%}$
Table 2: Comparison of Average Steps of all methods on the C-WAH task. ”Symbolic Obs” and ”Visual Obs” denote symbolic and visual observation settings, respectively. MHP uses a single agent while all others adopt two agents. indicates results quoted from CoELA. The subscript value like ${+33\%}$ in 75${}_{\ +33\%}$ denotes the Efficiency Improvement.
Baselines. We adopt three types of methods as our baseline: 1) classical agents, including MCTS-based Hierarchical Planner (MHP(Puig et al., 2020) and Rule-based Hierarchical Planner (RHP(Gan et al., 2022). 2) LLM-driven agents, including CoELA Zhang et al. (2023b), ProAgent (Zhang et al., 2023a), and RoCo (Mandi et al., 2023). CoELA Zhang et al. (2023b) features a modular framework for multi-agent planning, communication, and complete long-horizon tasks, but generate independent short-term plan for each agent. ProAgent (Zhang et al., 2023a), and RoCo (Mandi et al., 2023) generate joint plans for cooperative agents, and introduce a reflection loop or environment feedback for plan validation. 3) Heterogeneous agents, where agents with different planning and communication mechanisms collaborate within the same environment. See more details in Appendix A.1 and  A.2.